# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v1

@@ 5.2 Collaborative Semantic Segmentation
To further evaluate our approach, we focus on collaborative semantic segmentation by conducting experiments with real-world data from aerial-ground robots. We use the dataset CoPeD (Zhou et al., 2024), with one aerial robot and one ground robot, in two different real-world scenarios of the indoor NYUARPL and the outdoor HOUSEA. For more details about the dataset, please refer to (Zhou et al., 2024). Additionally, we introduce noise to the RGBD data collected by the ground robot. For both aerial and ground robots, RGB and depth data are used during training, while only RGB data is used during testing in CAML.
@@
We adopt the FCN (Long et al., 2015) architecture as the backbone for semantic segmentation. To process RGB and depth data locally for each robot, we use ResNet-18 (He et al., 2016) as the encoder to extract feature maps of size $7\times 7$. For more details about the experimental setup, please refer to the Appendix A.3.1. We first train a teacher model offline with aerial-ground robots collaboration using cross-entropy loss, where each robot has both RGB and depth data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for both aerial and ground robots through KD. The KD process is similar to that of the collaborative decision-making in CAV, but here we use a cross-entropy loss as the student task loss. For the detailed training settings, please see Appendix A.4.1.
We evaluate performance using the Mean Intersection over Union (mIoU) metric, which quantifies the average overlap between predicted segmentation outputs and ground truth across all classes. We compare the performance of CAML with other baselines including AML (Shen et al., 2023) and FCN (Long et al., 2015). In the AML approach, only the ground robot operates, with RGB and depth data available during training but only RGB data used for testing. The FCN approach involves only the ground robot operating with RGB data for both training and testing.
@@
We first present the experimental results of baselines comparison in Table 1, where CAML demonstrates superior performance in terms of mIoU across both indoor and outdoor environments. Specifically, CAML achieves an improvement of mIoU for ${\bf 8.88\%}$ in indoor scenario and ${\bf 10.61\%}$ in outdoor scenario compared to AML (Shen et al., 2023). We show the qualitative results in Fig. 8 in the Appendix A.3.2. Despite the noisy input image from the ground robot, CAML produces predictions that are closest to the ground truth. This improvement is attributed to CAML’s multi-agent collaboration, which provides complementary information to enhance data coverage and offers a more comprehensive understanding of the scenes. Additionally, the utilization of auxiliary depth data during training results in more precise segmentation outputs. We also investigate another variant of CAML, called Pre-fusion CAML, as ablation studies. Both CAML and Pre-fusion CAML have their advantages, and CAML can easily shift to Pre-fusion CAML because of the flexibility of our framework. Please refer to the Appendix A.3.3 for more details.
Table 1: Baseline Comparison of Semantic Segmentation on real-world dataset CoPeD (Zhou et al., 2024) using aerial-ground robots in indoor and outdoor environments. CAML achieves the highest mIoU in both environments, with upto 10.61% higher accuracy.
@@
FCN (Long et al., 2015) 51.20 56.22
AML (Shen et al., 2023) 55.89 60.32
CAML 60.05 66.83
Improvement over SOTA 4.16-8.88 6.51-10.61
@@ 6 Conclusions
In conclusion, we propose Collaborative Auxiliary Modality Learning (CAML), a unified framework for multi-agent multi-modality systems. Unlike prior methods that either focus on multi-agent collaboration without modality reduction or address multi-modality learning in single-agent settings, CAML integrates both aspects. It enables agents to collaborate using shared modalities during training while allowing efficient, modality-reduced inference. This not only lowers computational costs and data requirements at test time but also enhances predictive accuracy through multi-agent collaboration. We provide a theoretical analysis of CAML in terms of uncertainty reduction and data coverage, highlighting its advantages over AML. CAML demonstrates up to a $58.13\%$ improvement in accident detection for connected autonomous driving in complex scenarios and up to a $10.61\%$ mIoU gain in real-world aerial-ground collaborative semantic segmentation. These improvements underscore the practical implications of our framework. For limitations and future work, please see the Appendix A.6.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning, especially multi-agent multi-modality systems. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Appendix
A.1 Connected Autonomous Driving Scenarios
@@
We utilize a connected autonomous driving environment that integrates CARLA (Dosovitskiy et al., 2017) with AutoCast (Qiu et al., 2021). Our evaluation focuses on three complex and accident-prone traffic scenarios, characterized by limited sensor coverage or obstructed views. These scenarios are realistic and include background traffic of 30 vehicles. They involve challenging interactions such as overtaking, lane changing, and red-light violations, which inherently increase the risk of accidents: (1) Overtaking: A sedan is blocked by a truck on a narrow, two-way road with a dashed centerline. The truck also obscures the sedan’s view of oncoming traffic. The ego vehicle must decide when and how to safely pass the truck. (2) Left Turn: The ego vehicle attempts a left turn at a yield sign. Its view is partially blocked by a truck waiting in the opposite left-turn lane, reducing visibility of vehicles coming from the opposite direction. (3) Red Light Violation: As the ego vehicle crosses an intersection, another vehicle runs a red light. Due to nearby vehicles waiting to turn left, the ego vehicle’s sensors are unable to detect the violator.
@@
Figure 6: Three accident-prone scenarios in connected autonomous driving: overtaking, left turn, and red light violation.
A.2 Data Coverage
@@
We present a qualitative example highlighting how multi-agent collaboration provides complementary information to enhance data coverage. In a red-light violation scenario for connected autonomous driving, as shown in the following figure, the ego vehicle’s view is obstructed, rendering the occluded vehicle invisible. However, collaborative vehicles are able to detect the occluded vehicle, providing critical complementary information. This additional data helps the ego vehicle overcome its occluded view, enabling it to make more informed decisions and avoid potential collisions with the occluded vehicle.
@@ A.3.1 Experimental Setup
We resize the input RGB and depth images to $224\times 224$. To process RGB and depth data locally for each robot, we use ResNet-18 (He et al., 2016) as the encoder to extract feature maps of size $7\times 7$. The RGB features from both robots are shared and fused through channel-wise concatenation, and the depth features are processed similarly. Then we apply $1\times 1$ convolution to reduce the fused feature maps to the original channel dimensions for RGB and depth, respectively. We subsequently apply cross-attention to fuse the RGB and depth feature maps to generate multi-agent multi-modal feature aggregations. These aggregated features are passed through the decoder and upsampled to produce an output map matching the input image size.
A.3.2 Qualitative Results
We present the qualitative results of collaborative semantic segmentation using real-world data from aerial-ground robots in the following figure. Despite the noisy input image from the ground robot, CAML produces predictions closest to the ground truth. This performance is attributed to its multi-agent collaboration, which provides complementary information to enhance viewpoints, and its utilization of multi-modal depth data during training, enabling more precise segmentation outputs.
A.3.3 Ablation Studies
In the ablation studies, we explore another variant of CAML called Pre-fusion CAML, applied to the experiment of aerial-ground robots collaborative semantic segmentation. However, it is important to note that this variant can be applied to other domains and experiments as well. In this variant, each robot first locally extract feature maps of size $7\times 7$ for both RGB and depth modalities. Instead of separately fusing the RGB and depth features between the robots, we first fuse the feature maps of RGB and depth within each single robot using cross-attention. Then we share and merge the fused RGBD features between robots via concatenation. We also apply $1\times 1$ convolution to reduce the feature maps to the original channel dimensions. The multi-agent, multi-modal feature aggregations then pass through the decoder. Finally, we obtain the output map by upsampling to match the input image size. The mIoU of the Pre-fusion CAML is similar to that of CAML, achieving $59.16\%$ and $65.78\%$ for indoor and outdoor environments, respectively. By comparison, CAML achieves $60.05\%$ and $66.83\%$ in the same settings. Although the fusion order is different, both versions benefit from robust feature aggregation and multi-agent collaboration, which ultimately results in better segmentation performance.
@@ A.3.3 Ablation Studies
A.4 Complexity Analysis
A.4.1 Comparative Training Complexity
We report the training complexity of AML (Shen et al., 2023) and CAML for the experiments of collaborative decision-making in CAV, and collaborative semantic segmentation for aerial-ground robots in Table 2 and Table 3, respectively. For the experiments, we employ a batch size of 32 and the Adam optimizer (Kingma, 2014) with an initial learning rate of $1e{-3}$, and a Cosine Annealing Scheduler (Loshchilov & Hutter, 2016) to adjust the learning rate over time. The model is trained on an Nvidia RTX 3090 GPU with AMD Ryzen 9 5900 CPU and 32 GB RAM for 200 epochs.
@@ A.4.1 Comparative Training Complexity
A.4.2 Time and Space Complexity
In CAML, the agents’ embeddings are shared based on whether the system operates in a centralized or decentralized manner. If the system is a centralized, all collaborative agents share their data with one designated ego agent for centralized processing. Each of the $N-1$ collaborative agents performs its local computation independently, with a time complexity of $O(T_{c})$ and a space complexity of $O(S_{c})$, where $T_{c}$ represents the time required for local computation, and $S_{c}$ is the associated space. Thus, the total computation time and space complexities for all collaborative agents are $O(T_{c}(N-1))$ and $O(S_{c}(N-1))$, respectively. For simplicity, assuming each communication from one collaborative agent to the ego agent consumes $O(D)$ time complexity and $O(M)$ space complexity, where $D$ is the time required for communication and $M$ is the corresponding space. Therefore, the total communication time and space complexities for gathering information at the ego agent are $O(D(N-1))$ and $O(M(N-1))$, respectively. Then the ego agent aggregates the received data, running a model, having a time and space complexity $O(T_{e})$ and $O(S_{e})$, where $T_{e}$ and $S_{e}$ represent the time and space required for the ego agent’s computation. So the total time and space complexities are $O(T_{c}(N-1)+D(N-1)+T_{e})$ and $O(S_{c}(N-1)+M(N-1)+S_{e})$, respectively.
If the system is decentralized, each agent performs its local computation and shares information with other agents. For simplicity, let the local computation for a single agent has a time complexity of $O(T)$, where $T$ is the time required for local computation. Assume that communication from one agent to another agent requires $O(D)$ time complexity and $O(M)$ space complexity, where $D$ represents the time of communication between two agents, and $M$ denotes the space required for such communication. For $N$ agents, the total computation time complexity is $O(NT)$. In the worst case, each agent share data with all other agents, this can result in $O(N^{2}D)$ for pairwise sharing. So the total time complexity is $O(NT+N^{2}D)$. For space complexity, the storage requirement for all agents is $O(NS)$, where $S$ is the space needed per agent. Communication between agents adds an additional complexity of $O(N^{2}M)$. So the total space complexity is $O(NS+N^{2}M)$. In the typical case, if each agent communicates with only other $k$ agents ($k\ll N$) rather than all $N-1$ agents. The total time and space complexities become $O(NT+NkD)$ and $O(NS+NkM)$, respectively.
@@ A.5 Knowledge Distillation
We begin by training a teacher decision-making model $\mathcal{T}$ offline using both RGB and LiDAR data, with a binary cross-entropy loss: $\mathcal{L}_{BCE}(y,\mathcal{T})=-\mathbb{E}_{\mathcal{D}}\big{[}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\big{]}$, where $\mathcal{D}$ is the dataset, $y_{i}$ is the ground truth indicating whether the vehicle should brake, $p_{i}$ is the predicted probability by the teacher model $\mathcal{T}$. The student model $\mathcal{S}$ is trained to mimic the behavior of the teacher model while having less modalities. For each data point, the student model receives the same RGB image that the teacher model was given. The loss for the student model is a combination of two terms: the distillation loss using KL divergence between the student output and teacher output (soft targets), and the student task loss, which is the binary cross entropy loss between the student output and the true labels (hard targets). The soft targets from the teacher enrich learning with class similarities, while hard targets ensure alignment with true labels. The soft targets are generated by applying a temperature scaling to the logits. The scaled logits are defined as: $z_{i}=\frac{\exp{(z_{i}/t)}}{\exp{(z_{0}/t)}+\exp{(z_{1}/t)}}$, where $z_{i}$ is the logit for class $i$ and $t=3.0$ is the temperature parameter. The distillation loss is defined as: $\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})=-\sum z_{i}^{\mathcal{T}}\log(z_{i}^{\mathcal{S}})$, where $z_{i}^{\mathcal{T}}$ and $z_{i}^{\mathcal{S}}$ are the soft target probability from the teacher and student model, respectively. The overall loss for the student model is a weighted sum of the distillation loss and the binary cross-entropy loss: $\mathcal{L}_{\mathcal{S}}=(1-\alpha)\mathcal{L}_{BCE}(y,\mathcal{S})+\alpha t^{2}\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})$, where $\alpha=0.5$ controls the trade-off between the two losses. After the training of knowledge distillation process, we obtain a student model that uses only RGB data while learning from a teacher model that has access to both RGB and LiDAR data. This enables the student model to be effective during testing with only RGB data. Additionally, by leveraging knowledge distillation, the student model benefits from the additional insights provided by the LiDAR data during training, learning more effectively compared to training solely with RGB data.

### v2

@@ 4.2 Collaborative Semantic Segmentation
To further evaluate our approach, we focus on collaborative semantic segmentation by conducting experiments with real-world data from aerial-ground robots. We use the dataset CoPeD [41], with one aerial robot and one ground robot, in two different real-world scenarios of the indoor NYUARPL and the outdoor HOUSEA. For more details about the dataset, please refer to [41]. Additionally, we introduce noise to the RGBD data collected by the ground robot. For both aerial and ground robots, RGB and depth data are used during training, while only RGB data is used during testing in CAML.
@@
We adopt the FCN [23] architecture as the backbone for semantic segmentation. To process RGB and depth data locally for each robot, we use ResNet-18 [14] as the encoder to extract feature maps of size $7\times 7$. Please refer to Appendix A.3.1 for more details of the experimental setup. We first train a teacher model offline with aerial-ground robots collaboration using cross-entropy loss, where each robot has both RGB and depth data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for both aerial and ground robots through KD. The KD process is similar to that of the collaborative decision-making in CAV, but here we use a cross-entropy loss as the student task loss. Please see Appendix A.4 for the detailed training settings. We evaluate performance using the Mean Intersection over Union (mIoU) metric, which quantifies the average overlap between predicted segmentation outputs and ground truth across all classes. We compare the performance of CAML with other baselines including AML [32] and FCN [23]. In the AML approach, only the ground robot operates, with RGB and depth data available during training but only RGB data used for testing. The FCN approach involves only the ground robot operating with RGB data for both training and testing.
@@
We first present the experimental results of baselines comparison in Table 3, where CAML demonstrates superior performance in terms of mIoU across both indoor and outdoor environments. Specifically, CAML achieves an improvement of mIoU for ${\bf 8.9\%}$ in indoor scenario and ${\bf 10.6\%}$ in outdoor scenario compared to AML [32]. We also show the qualitative results in Fig. 5. As we can see, despite the noisy input image from the ground robot, CAML produces predictions that are closest to the ground truth. This improvement is attributed to CAML’s multi-agent collaboration, which provides complementary information to enhance data coverage and offers a more comprehensive understanding of the scenes. Additionally, the utilization of auxiliary depth data during training results in more precise segmentation outputs.
Table 3: Baseline Comparison of Semantic Segmentation on real-world dataset CoPeD [41] using aerial-ground robots in indoor and outdoor environments. CAML achieves the highest mIoU in both environments, with upto 10.6% higher accuracy.
@@
FCN [23] 51.2 56.2
AML [32] 55.9 60.3
CAML 60.1 66.8
Improvement over SOTA 4.2-8.9 6.5-10.6
Ablation Studies.
We also investigate another variant of CAML, called Pre-fusion CAML, as ablation studies. In this variant, each robot first locally extracts feature maps of size $7\times 7$ for both RGB and depth modalities. Instead of separately fusing the RGB and depth features between the robots, we first fuse the feature maps of RGB and depth within each single robot using cross-attention. Then we share and merge the fused RGBD features between robots via concatenation. We also apply $1\times 1$ convolution to reduce the feature maps to the original channel dimensions. The multi-modal multi-agent feature aggregations then pass through the decoder. Finally, we obtain the output map by upsampling to match the input image size. The mIoU of the Pre-fusion CAML is similar to that of CAML, achieving $59.2\%$ and $65.8\%$ for indoor and outdoor environments, respectively. Although the fusion order is different, both versions benefit from robust feature aggregation and multi-agent collaboration, which ultimately results in better segmentation performance. And CAML can easily shift to Pre-fusion CAML because of the flexibility of our framework. Please refer to Appendix A.3.2 for more details.
@@ 5 Conclusions
In conclusion, we propose Collaborative Auxiliary Modality Learning (CAML), a unified framework for multi-modal multi-agent systems. Unlike prior methods that either focus on multi-agent collaboration without modality reduction or address multi-modal learning in single-agent settings, CAML integrates both aspects. It enables agents to collaborate using shared modalities during training while allowing efficient, modality-reduced inference. This not only lowers computational costs and data requirements at test time but also enhances predictive accuracy through multi-agent collaboration. We provide an intuitive analysis of CAML in terms of data coverage and information gain, justifying its effectiveness. CAML demonstrates up to a $58.1\%$ improvement in accident detection for connected autonomous driving in complex scenarios and up to a $10.6\%$ mIoU gain in real-world aerial-ground collaborative semantic segmentation. These improvements underscore the practical implications of our framework, especially for resource-constrained environments. For limitations and future work, please see Appendix A.6.
@@ Appendix A Appendix
A.1 Analysis
We provide an intuitive analysis to justify the effectiveness of CAML and illustrate why it outperforms single-agent approaches such as AML. Please note that this analysis is intended to offer intuitive insights and explanations into the benefits of CAML, rather than serving as a formal mathematical proof, as this may be intractable in this context.
We analyze from two key perspectives: Data Coverage and Information Gain. We aim to address the following key questions: (a) Does the collaboration of multiple agents provide complementary information that increases data coverage? Specifically, does combining observations from each agent lead to a more accurate and comprehensive prediction compared to using a single agent? (b) Does the collaboration increase the mutual information between the observations and the true label?
Data Coverage.
To address Question (a), we study data coverage and information provided by each agent in a multi-agent system. Let the entire data space be denoted as ${\mathcal{D}}$, which consists of various subsets. Each agent ${\mathcal{A}}_{i}$ in the system covers a subset of this data space: ${\mathcal{C}}_{i}\subseteq{\mathcal{D}}$. The overall coverage by the system is given by the union of all subsets covered by individual agents: ${\mathcal{C}}_{multi}=\cup_{i=1}^{N}{\mathcal{C}}_{i}$. This ensures that $|{\mathcal{C}}_{multi}|\geq\max|{\mathcal{C}}_{i}|$. If only a single agent is available, it can only observe a portion of the data space, leaving parts of the space unobserved, which leads to incomplete information for estimating the true label $y$. We show a qualitative example of multi-agent collaboration providing complementary information to enhance data coverage in Fig. 7 in Appendix A.2.2.
From a probabilistic perspective, when multi-agent collaboration is in place, the combined likelihood $P(X|y)$ is modeled as a multivariate distribution. This approach provides a broader and more accurate representation of the data space by integrating information from all agents and modeling the dependencies and correlations between them. Compared to a univariate distribution $P(x_{i}|y)$ for a single agent ${\mathcal{A}}_{i}$, the multivariate distribution covers a larger portion of the data space ${\mathcal{D}}$, thus enhancing data coverage. This allows the exploration of more complex patterns, relationships, and complementary information from different agents. By capturing a richer set of interactions and correlations among the agents’ observations, the multivariate distribution supports more informed decision-making. The model’s predictions are based on a comprehensive view of the environment, thus leading to more accurate outcomes.
Information Gain.
To address Question (b) about information gain, we analyze using information theory. Let $I(y;x_{i})$ represent the mutual information between the true label $y$ and agent ${\mathcal{A}}_{i}$’s observation $x_{i}$, which quantifies how much information $x_{i}$ provides about the estimation of $y$. The joint mutual information between $y$ and the set of all observations $X$ is $I(y;X)$. In the context of multi-agent collaboration, the joint observations $X$ from multiple agents typically provide more comprehensive information about the true label $y$ compared to the observation of any single agent. Therefore, the mutual information $I(y;X)$ is always greater than or equal to the mutual information from a single agent: $I(y;X)\geq I(y;x_{i})$. We can formally justify this using the chain rule for mutual information:
$I(y;X)$ $=I(y;x_{1},\dots,x_{n})$ (1)
$=I(y;x_{1})+I(y;x_{2}\mid x_{1})+\cdots+I(y;x_{n}\mid x_{1},\dots,x_{n-1}).$
Each term ($I(y;x_{k}\mid x_{1},\dots,x_{k-1})\geq 0$) is non-negative because mutual information is always non-negative. It follows that:
$I(y;X)\geq I(y;x_{i})\quad\text{for any individual }x_{i}.$ (2)
Adding more observations (e.g., expanding from $x_{i}$ to $X$) does not reduce the information about $y$. Even if some observations are redundant (e.g., $x_{j}$ duplicates $x_{i}$), the joint mutual information $I(y;X)$ remains at least as large as $I(y;x_{i})$. Thus, the combined observations from multi-agent collaboration generally provide more information about $y$ than a single observation, improving the overall estimate.
A.2 Connected Autonomous Driving
A.2.1 Dataset Details
@@
The dataset is generated using the AUTOCASTSIM [29] benchmark, which features three complex and accident-prone traffic scenarios for connected autonomous driving, characterized by limited sensor coverage or obstructed views. These scenarios are realistic and include background traffic of 30 vehicles. They involve challenging interactions such as overtaking, lane changing, and red-light violations, which inherently increase the risk of accidents: (1) Overtaking: A sedan is blocked by a truck on a narrow, two-way road with a dashed centerline. The truck also obscures the sedan’s view of oncoming traffic. The ego vehicle must decide when and how to safely pass the truck. (2) Left Turn: The ego vehicle attempts a left turn at a yield sign. Its view is partially blocked by a truck waiting in the opposite left-turn lane, reducing visibility of vehicles coming from the opposite direction. (3) Red Light Violation: As the ego vehicle crosses an intersection, another vehicle runs a red light. Due to nearby vehicles waiting to turn left, the ego vehicle’s sensors are unable to detect the violator.
Following the setup established by prior works COOPERNAUT [5] and STGN [8], we collect 24 data trails for each scenario, using 12 trails for training and the remaining 12 for testing. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. These variations include different types of vehicles present in the scene, varying initial positions and trajectories of collaborative vehicles, difference in traffic flow and interactions, resulting in different visual, LiDAR inputs and control actions for each trail. Each trail contains RGB images and LiDAR point clouds from multiple vehicles, resulting in a substantially large dataset with a total size of approximately 400 GB.
@@
Figure 6: Three accident-prone scenarios in connected autonomous driving benchmark AUTOCASTSIM: overtaking, left turn, and red light violation.
A.2.2 Data Coverage
@@
We present a qualitative example highlighting how multi-agent collaboration provides complementary information to enhance data coverage in Fig. 7. In a red-light violation scenario for connected autonomous driving, as shown in the following figure, the ego vehicle’s view is obstructed, rendering the occluded vehicle invisible. However, collaborative vehicles are able to detect the occluded vehicle, providing critical complementary information. This additional data helps the ego vehicle overcome its occluded view, enabling it to make more informed decisions and avoid potential collisions with the occluded vehicle.
@@
A.2.3 Behavior Cloning Details
In the CAV task, we apply Behavior Cloning (BC) following prior works COOPERNAUT [5] and STGN [8]. The process involves first running an expert agent to collect data from both the ego vehicle and collaborative vehicles. This data includes RGB images, LiDAR point clouds, and the control actions of the ego vehicle, specifically, braking decisions. Please see more details of the data collection in Appendix A.2.1.
To integrate BC into our multi-agent setup, we train a model that takes as input the RGB and LiDAR data from both the ego and collaborative vehicles, and it outputs the control actions for the ego vehicle. BC is trained using a cross-entropy loss to detect accident-prone cases. The purpose of BC is to let the ego vehicle effectively mimic the expert’s behavior, allowing it to better assess whether a situation is dangerous or not and make informed decisions, such as braking or continuing to drive. This enhances safety in accident-prone environments by ensuring more accurate and context-aware decision-making.
@@ A.3.1 Experimental Setup
We resize the input RGB and depth images to $224\times 224$. To process RGB and depth data locally for each robot, we use ResNet-18 [14] as the encoder to extract feature maps of size $7\times 7$. The RGB features from both robots are shared and fused through channel-wise concatenation, and the depth features are processed similarly. Then we apply $1\times 1$ convolution to reduce the fused feature maps to the original channel dimensions for RGB and depth, respectively. We subsequently apply cross-attention to fuse the RGB and depth feature maps to generate multi-agent multi-modal feature aggregations. These aggregated features are passed through the decoder and upsampled to produce an output map matching the input image size.
A.3.2 Advantages of CAML and Pre-fusion CAML
@@
A.4 Training Complexity
We report the training complexity of AML [32] and CAML for the experiments of collaborative decision-making in CAV, and collaborative semantic segmentation for aerial-ground robots in Table 4 and Table 5, respectively. For the experiments, we employ a batch size of 32 and the Adam optimizer [18] with an initial learning rate of $1e{-3}$, and a Cosine Annealing Scheduler [24] to adjust the learning rate over time. The model is trained on an Nvidia RTX 3090 GPU with AMD Ryzen 9 5900 CPU and 32 GB RAM for 200 epochs.
@@ A.5 Knowledge Distillation
We begin by training a teacher decision-making model $\mathcal{T}$ offline using both RGB and LiDAR data, with a binary cross-entropy loss: $\mathcal{L}_{BCE}(y,\mathcal{T})=-\mathbb{E}_{\mathcal{D}}\big{[}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\big{]}$, where $\mathcal{D}$ is the dataset, $y_{i}$ is the ground truth indicating whether the vehicle should brake, $p_{i}$ is the predicted probability by the teacher model $\mathcal{T}$. The student model $\mathcal{S}$ is trained to mimic the behavior of the teacher model while having less modalities. For each data point, the student model receives the same RGB image that the teacher model was given. The loss for the student model is a combination of two terms: the distillation loss using KL divergence between the student output and teacher output (soft targets), and the student task loss, which is the binary cross entropy loss between the student output and the true labels (hard targets). The soft targets from the teacher enrich learning with class similarities, while hard targets ensure alignment with true labels. The soft targets are generated by applying a temperature scaling to the logits. Following prior works [15, 32, 34], the scaled logits are defined as: $z_{i}=\frac{\exp{(z_{i}/t)}}{\exp{(z_{0}/t)}+\exp{(z_{1}/t)}}$, where $z_{i}$ is the logit for class $i$ and $t=4.0$ is the temperature. The distillation loss is defined as: $\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})=-\sum z_{i}^{\mathcal{T}}\log(z_{i}^{\mathcal{S}})$, where $z_{i}^{\mathcal{T}}$ and $z_{i}^{\mathcal{S}}$ are the soft target probability from the teacher and student model, respectively. The overall loss for the student model is a weighted sum of the distillation loss and the binary cross-entropy loss: $\mathcal{L}_{\mathcal{S}}=(1-\alpha)\mathcal{L}_{BCE}(y,\mathcal{S})+\alpha t^{2}\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})$ with $\alpha=0.9$ as the weight. After the training of knowledge distillation process, we obtain a student model that uses only RGB data while learning from a teacher model that has access to both RGB and LiDAR data. This enables the student model to be effective during testing with only RGB data. Additionally, by leveraging knowledge distillation, the student model benefits from the additional insights provided by the LiDAR data during training, learning more effectively compared to training solely with RGB data.
Training a teacher model and then distilling it into a student model allows the student model benefit from the richer information that the teacher model has. Since student models are typically smaller or operate with fewer modalities, they are designed to be more efficient, with fewer parameters or reduced complexity. Deploying such a student model in a resource-constrained environment is crucial, especially when the teacher model is too costly to deploy due to its computation requirements.